# AI-Based Stroke Rehabilitation Domiciliary Assessment System with ST-GCN Attention

### v1

@@ VI Conclusion
Our approach offers the advantage of providing assessment scores without requiring a control group of healthy reference motions. Collecting rehabilitation movement datasets is costly regardless of whether the subjects are able-bodied or impaired. Moreover, as discussed earlier, rehabilitation assessment is conducted in terms of achieving functional activity and participation. Thus, the significance of our method lies in performing patient-centered evaluation, focusing on actual patient movements and analyzing their performance without the need for non-patient reference data. Nevertheless, the present model remains limited to assessment and does not define or prescribe recovery strategies within a comprehensive home-based rehabilitation system. Future research will focus on developing an integrated higher-level framework capable of generating recovery plans, based on more advanced analyses of individual rehabilitation exercises.

### v2

@@ VI Conclusion
Our approach provides assessment scores without requiring a non-disabled control group or reference motions. Collecting rehabilitation movement data is costly for both impaired and non-impaired participants. Moreover, rehabilitation assessment focuses on functional activity and participation. Accordingly, our method supports patient-centered evaluation by analyzing each patient’s own movements and performance without relying on non-patient reference data.
Nevertheless, the proposed model (RAST-G@) is limited to assessment and does not yet address treatment planning within a comprehensive home-based rehabilitation system. Future work will focus on an integrated higher-level framework that can generate personalized recovery plans based on more detailed analyses of individual exercises.